OpenStudy (anonymous):

A number is called "even-odd" if it is halfway between n even integer and an odd integer. If x is an even-odd number. Which of the following must be true? a) 2x is an integer b) 2x is even-odd c) x is halfway between two even integers

OpenStudy (tkhunny):

Here's your 'even-odd' numbers, \(\dfrac{2k + (2j+1)}{2}\), where j and k are integers. Now what?

OpenStudy (anonymous):

2x will always be odd

OpenStudy (anonymous):

\[\frac{2k+2j+1}{2}=x\\2x=2(k+j)+1\]

OpenStudy (anonymous):

i wrote the question wrong

OpenStudy (anonymous):

A number is called "even-odd" if it is halfway between an even integer and an odd integer. If x is an even-odd number. Which of the following must be true? a) 2x is an integer b) 2x is even-odd c) x is halfway between two even integers

OpenStudy (anonymous):

the answer is a

OpenStudy (anonymous):

and in particular, an odd integer

OpenStudy (anonymous):

ok